OpenStudy (anonymous):

What is the surface area of the cylinder? Diameter of the circle=12 Height of cylinder=15 I keep getting 791 and its not an option.

OpenStudy (anonymous):

WAIT, derp I am stupid, friggin pie is part of the answer, let me just go cry in a corner

OpenStudy (lgbasallote):

surface area = \(2\pi r^2 h + 2\pi r h\) since diameter = 12 then that means radius is 6 \[SA = 2 \pi (6)^2 (15) + 2 \pi (6)(15)\] \[SA = 2 \pi (36)(15) + 2 \pi (90)\] is this what you did?

OpenStudy (lgbasallote):

oh lol

OpenStudy (anonymous):

Yeah, I was doing it the more difficult way I guess, thanks for helping though!

OpenStudy (lgbasallote):

<tips hat>

OpenStudy (anonymous):

Wait, @lgbasallote I think the formula is \(2πr^{2} + 2πrh\)

OpenStudy (lgbasallote):

oops lol yeah delete 15 there *facepalm*